OpenStudy (anonymous):

Am I doing this right so far? What is the 15th term of the arithmetic sequence 16, 23, 30, ? First term: 16 Common ratio: 23/16 = 1.4375

hartnn (hartnn):

why common ratio ? thats for geometric sequence, we are dealing with arithmetic sequence here.

hartnn (hartnn):

so, find common difference d=.... ?

OpenStudy (anonymous):

oh my error

OpenStudy (anonymous):

oh right!!! ok let me seee

OpenStudy (anonymous):

so i find a number that goes into all of the numbers above?

hartnn (hartnn):

you find a number which when added to current term, gives you next term

hartnn (hartnn):

what would you add to 16 to get 23 what would you add to 23 to get 30 ?

OpenStudy (anonymous):

16 + 7?

OpenStudy (anonymous):

omg it's 7!!! :))

hartnn (hartnn):

thats correct, so common difference d=30-23=23-16 =7 now you know the formula for n'th term of arithmetic sequence ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, what you got 15th terms as ?

OpenStudy (anonymous):